Saturday, Mar. 05: DAY Shift
From Hall A Wiki
Go back to Daily Run Plan Daily Run Plans
- Production data (Kin48_3)
- Continue production mode ***
last update Feb 29
- Polarized beam at 15 uA on LH2 15 cm through Compton chicane !
- DVCS is running kin_48_3 [here], as off 02-29 10 am we have accumulate 1 C out of our goal of 4.7 C.
- LHRS is partially detuned: [here]
- LHRS=2.92GeV, Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.
- beam locks: no position lock, energy lock in Arc 2: [here]
- the straw chambers in the LHRS are off
- DVCS daq configuration is [this one], we are measuring about 2% DT at 15uA.
- Take Compton run every 2 h
- Dealing with Issues:
- Can't control the Alpha Y axis of the raster. (MCC tried reboot but failed) - 1AM
- I came to CH and called to Bill to check
- Bill Gunning came to replace the PS (the escorted access)
- New PS initialization failed / dump diffuser issue, MCC reboot again but IOChla fail again (multi-times escorted access)
- From MCC(9:00AM) meeting, PD said it might be a remote reset switch (needs an analysis by expert).
- If happen again, the remote reset switch will be replaced and more investigate detail during access.
- Raster seemed controllable by MCC (look at the Oscilloscope), but spot++ data showed X did not change at all any current
- Expert called, Bob came and investigating... HALOG
- Meantime, CHL is crashed. Hall-A is now controlled access - 1PM
- Repair raster is done ! HALOG
- After beam lost for considerable time (like beam study, pass change,...)
- Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++
Please check more info on the white board in the Counting House ~!!